# An Overview of Cyber Security Funding for Open Source Software

### v1

@@ 5 Results
The dissemination of the qualitative results can be started by noting that ten out of the $68$ funding cases did not address cyber security. These cases are thus excluded from the answers given to RQ.4. Then, regarding RQ.1, a concise thematic summary is presented in Table 1 about the types of OSS funded by the two funding bodies.
As can be seen, open source software supply chains, including package managers for installing, updating, and maintaining dependencies, have been the most frequently funded target. The result is not surprising because the security issues with OSS supply chains are well-recognized (Ohm et al.,, 2020; Ruohonen et al.,, 2021; Zimmermann et al.,, 2019). These issues affect large user and deployment bases, including in the commercial software industry due to the widespread use of OSS components and dependencies. Supply chain security and management are also more generally seen as being among the foremost difficult global issues in the increasingly globalized and digitalized world (Wu,, 2024). Thus, it is no wonder that the STA and OpenSSF too have recognized their criticality.
@@
Network libraries and protocol implementations have also been frequently funded particularly by the STA. As could be expected, DNS and BGP implementations and their projects are among the funding recipients. Also other networking libraries and implementations crucial for the whole Internet are among the recipients, including those implementing the transport layer security (TLS) and secure shell (SSH) cryptographic protocols (OpenSSF, 2024c, ; STA, 2024d, ). It is worth mentioning also funding for a project implementing end-to-end encryption protocols as well as four separate funding decisions for projects implementing the pretty good privacy (PGP) cryptography protocol. Of the newer networking protocols related to security, WireGuard has also received funding.
@@
Finally, regarding the concept of criticality, it is understandable that also operating systems and their low-level software components have received funding. An example would be FreeBSD, but the OpenSSF has also granted funding to the Linux kernel, which is closely associated with the OpenSSF’s parent organization, the Linux Foundation. Although the data is too limited for answering to the associated why-question, this observation is enough to remark that the OSS funding bodies—like many but not necessarily all funding bodies and financing instruments in general, may not be entirely immune to potential biases and conflicts of interests. A further point is that criticality as a concept seems to align with the STA, 2024g’s () notion of foundational technologies without which the Internet may function poorly and the creation of new software may become difficult. The frequent funding of OSS supply chains further aligns with the OpenSSF, 2024d’s () emphasis of dependencies for defining criticality. However, there is also some room for criticism regarding the concept of criticality and its relation to the funding decisions. For instance, STA, 2024e has granted $450$ thousand euros for a project seeking to improve documentation about a JavaScript library. While documentation is important, it remains debatable how critical the library is with respect to the foundational technologies emphasized by the STA.
@@
Turning to RQ.2, the monetary amounts granted are visualized in Fig. 1. (It can be noted that at the time of writing the exchange rate between the EU’s euro and the United States dollar was almost even, and thus no conversion was needed.) The average is a little below four thousand euros, but there are a few outlying projects that have received even over a million euros. Without prior work, it is rather difficult to contemplate how large or small these amounts are. Nevertheless, consider an imaginary example of a professional software engineer earning roughly around five thousand euros a month; then, the average funding amount would already account for about six full-time software engineers working with a project for a year. This example allows to argue that the funding amounts are decent enough in the OSS domain. Although the OpenSSF does not provide data on the funding durations, most (84%) of the STA’s funding decisions have been for two years. This observation further reinforces the point about the interpretation of the money involved.
@@
The types of cyber security funded (RQ.4) summarized in Table 2 expectedly align closely with the earlier thematic summary in Table 1. In other words, supply chain security has been the most typical thematic cyber security area funded. In this regard, it is important to further note that some of the OSS projects funded have also addressed the design and implementation of so-called software bills of materials (SBOMs). These materials for keeping track of dependencies and other components are noteworthy for two reasons, the first being that SBOMs are also under active research in software engineering (Bi et al.,, 2024; Xia et al.,, 2023, among others). The second reason is that SBOMs too are present in the EU’s CRA, which mandates commercial software vendors to implement SBOMs for their products and deliver these to market surveillance authorities upon request. Although OSS projects and their so-called stewards are excluded from this regulatory requirement, the integration of OSS components into commercial software products might partially explain also the funding decisions for SBOMs by the STA and OpenSSF. Another partial explanation might be that also the funding bodies perceive SBOMs as important for improving supply chain security and software security of OSS in general.
@@
Then, regarding RQ.5, already the summary in Table 2 is enough to conclude that there is a divergence between the funding bodies also with respect to the types of cyber security funded. In particular, the STA has funded also many projects addressing encryption and cryptography, network security, and security testing, whereas the OpenSSF has been particularly active in funding of security audits, including those done by third parties. Having said that, there are also some projects that have received funding from both OSS funding bodies. The examples include FreeBSD and the Python Software Foundation, which is maintaining the popular and large repository for distributing software packages written in Python. The data is again too limited for far-reaching speculations, but, nevertheless, an argument can be raised that the two funding bodies might benefit from joint coordination. This point can be strengthened by noting that the data contains also some cases indicating joint ventures between different funds (OpenSSF, 2024c, ). Against this backdrop, it might be even reasonable to argue that the whole OSS funding landscape might benefit from a high-level coordination body through which all funds could collaborate.
Audits and auditing are important to underline because these have involved also commercial cyber security companies doing the auditing. Thus, the actual money involved may sometimes go to the private sector even though the goal of funding remains the same. This point deserves three additional brief remarks. The first remark is that the involvement of the private sector aligns the OSS funding topic toward so-called public-private partnerships (PPPs), which have been common in the cyber security domain in general (Ruohonen,, 2020). Insofar as OSS is interpreted to be on the public side, which would seem a reasonable interpretation with respect to the concept of public goods underneath the tragedy of the commons, PPPs are particularly visible in the OpenSSF case because the associated fund was established by donations from large technology companies, among them Amazon, Google, and Microsoft (OpenSSF,, 2023). The second remark follows: PPPs are typically also required for funding of large-scale applied scientific projects, including those funded by the EU’s Horizon Europe program. These two remarks logically lead to the final remark: the OSS funding topic is also about the so-called technology transfer model, which has been endorsed and debated in software engineering research for a long time (Mikkonen et al.,, 2018; Wohlin et al.,, 2000, among many others). Regarding the transfers involved, however, scientific research and academia are missing. Given the Germany’s involvement behind the STA, the transfers are primarily between the public sector, the public (OSS) domain, and the private sector.
Therefore, it could be contemplated whether adding academic software engineering research to the mixture would improve the security of OSS even further, perhaps opening also new collaborative avenues and further funding possibilities. There are also some indications in the data that some projects have already justified their funding applications with references to scientific research projects and the EU’s own software development projects (STA, 2024c, ). Thus, the earlier point about coordination can be extended toward a larger governance space, which, in Europe, involves also the cyber security regulations mentioned.
@@
  • 1.
    1 Excluding software bills of materials (SBOMs).
  • 2.
    2 Application programming interfaces (APIs).
@@
With respect to RQ.3, the funding rationale themes are summarized in Table 3. To put aside cyber security itself, which has sometimes been justified with explicit mentions of critical infrastructure (STA, 2024b, ), also the sustainability theme is strongly present as a rationale. Sustainability as a theme also aligns with the rather frequent rationales of software testing and plain old bug fixing. However, it is worthwhile to mention that some projects have approached the sustainability conundrum through identifying and then deprecating components and dependencies that have reached their end-of-life software life cycle stages or became technologically risky (STA, 2024c, ; STA, 2024d, ). In a similar vein, the earlier point about technological debt is visible in the few projects that have approached sustainability problems through wholesale rewrites and re-implementations. Such solutions are often interpreted as a sign of technical debt of the most severe kind (Verdecchia et al.,, 2021). However, the rewrites and re-implementations have not been only about technological debt and sustainability in general.
@@
The reason is that some of these have involved also transformations to other programming languages; hence, the rewrites and re-implementations also align with the memory safety theme in the earlier Table 2. The Rust programming language is the prime example in this regard. The language and its ecosystem have also received funding from both bodies. From the STA, 2024f the Rust project received funding for implementing SBOMs, while from the OpenSSF, 2024b it received funding for a myriad of security themes, including hiring a dedicated security engineer, audits, documentation, advocacy, and education, as well as new security tools and an establishment of a yet another grant program for mentoring and education. The Rust’s own grant program yet again highlights the potential benefits from better coordination. The OpenSSF, 2024b’s () funding case also indicates that the funding applications and their evaluations might benefit from a more structured approach. For instance, like in large-scale scientific grant applications, it could be contemplated whether structuring the themes and rationales into work packages would bring more structure and coherence. A further point is that the education, outreach, and documentation involved, and the accompanying overall theme in Table 3, again frames the OSS funding topic also toward academia. In this regard, it could be argued that the security themes funded by the STA and OpenSSF are also something that should be taught in higher education. The Rust case also demonstrates that there might be a need in the nearby future to update curricula with respect to programming languages.
@@
Finally, it is important to emphasize that the funding rationales have involved also implementing new features explicitly or implicitly related to software security and cyber security in general. In addition, also many other general software engineering themes are present, including performance and scalability, interoperability, reliability, usability, and software engineering processes and practices. These rationales allow to end the qualitative exposition by arguing that the tragedy of the (digital) commons alone can neither explain nor elaborate the whole topic of OSS funding. Also many other explanations are presumably present. The PPP aspect noted is among these explanations; the private sector also greatly benefits from OSS. Thus, further research is needed to better understand the business aspects behind OSS funding. Concepts such as so-called coopetition (Duc et al.,, 2017) might provide a theoretical path to follow in this regard.
6 Conclusion and Discussion
@@
However, cyber security and sustainability alone cannot fully explain the rationales for funding; therefore, neither is the tragedy of the (digital) commons sufficient as an overall theoretical explanation. In addition to raising this important argument, the paper made a notable contribution by linking together the research branches addressing the sustainability of OSS projects and critical infrastructure. Regarding the latter research branch, software’s role has often been downplayed to some extent, and the role of OSS has been almost nonexistent in existing research. A further contribution was made by elaborating how the OSS cyber security funding theme examined is related to regulations. Particularly the EU’s new CRA regulation is relevant also for open source software projects.
@@
The concept of criticality is the keyword behind the funding decisions. However, as was discussed, it is not entirely clear how the two funding bodies have used the concept in decision-making about funding applications. For instance, the corresponding working group established by OpenSSF, 2024d has emphasized dependencies, popularity, and related concepts in determining a project’s criticality. However, dependencies and popularity may align poorly with the STA’s notions about open digital infrastructures and foundational technologies. Consider DNS and BGP implementations as an example; they are clearly critical for the Internet’s functioning, but they may not be as popular as some other OSS projects with large user bases. Nor do they garner a large amount of dependencies. Thus, further research is required about the criticality concept and its practical use in funding decisions.
@@
In addition, an important point regarding further research is the evaluation of not only funding applications but the successes or failures of the projects funded. In this regard, OpenSSF, (2023) has noted that their evaluation has been “largely anecdotal” and that further goals include “highlighting the success factors to future grant recipients and using them in our grant review process”. The STA, 2024a’s () evaluation criteria for funding applications are a little more systematic; prevalence, relevance, sustainability, public interests, activities planned and their feasibility, and expertise of the applicants are evaluated. However, the Agency’s online materials do not say anything about whether funded OSS projects are evaluated post hoc after their completions.
@@
To this end, different evaluation criteria have been considered and compared also in existing research (Osborne et al.,, 2024), but it remains unclear how applicable they are in practice. Science funding provides a good example. On one hand, learning from the funding of scientific projects might help with the goal of improving evaluations. Provided that the point about potential biases and conflicts of interests is true even to some extent, the learning might offer means to improve objectivity in reviewing because objectivity has traditionally been highly valued in science. For instance, it is unclear whether the STA and OpenSSF have used external reviewers, and whether such reviewers would be seen as preferable. On the other hand, it could be also argued that science funding might be a poor reference because it is often rather bureaucratic.

### v2

@@ 5 Results
The dissemination of the qualitative results can be started by noting that ten out of the $68$ funding cases did not address cyber security. A good example of non-security funding would be a grant for the GNOME desktop environment (STA, 2024c, ). These cases are thus excluded from the answers given to RQ.4. Then, regarding RQ.1, a concise thematic summary is presented in Table 1 about the types of OSS funded by the two funding bodies.
As can be seen, open source software supply chains, including package managers for installing, updating, and maintaining dependencies, have been the most frequently funded target. The result is not surprising because the security issues with OSS supply chains are well-recognized (Hamer et al.,, 2025; Ohm et al.,, 2020; Ruohonen and Saddiqa,, 2025; Zimmermann et al.,, 2019). These issues affect large user and deployment bases, including in the commercial software industry due to the widespread use of OSS components and dependencies. Supply chain security and management are also more generally seen as being among the foremost difficult global issues in the increasingly globalized and digitalized world (Wu,, 2024). Thus, it is no wonder that the STA and OpenSSF too have recognized their criticality.
@@
Network libraries and protocol implementations have also been frequently funded particularly by the STA. As could be expected, DNS and BGP implementations and their projects are among the funding recipients. Also other networking libraries and implementations crucial for the whole Internet are among the recipients, including those implementing the transport layer security (TLS) and secure shell (SSH) cryptographic protocols (OpenSSF, 2024d, ; STA, 2024f, ). It is worth mentioning also funding for a project implementing end-to-end encryption protocols as well as four separate funding decisions for projects implementing the pretty good privacy (PGP) cryptography protocol. Of the newer networking protocols related to security, WireGuard has also received funding.
@@
Finally, regarding the concept of criticality, it is understandable that also operating systems and their low-level software components have received funding. An example would be FreeBSD (OpenSSF, 2024b, ), but the OpenSSF has also granted funding to the Linux kernel, which is closely associated with the OpenSSF’s parent organization, the Linux Foundation. Although the data is too limited for answering to the associated why-question, this observation is enough to remark that the OSS funding bodies—like many but not necessarily all funding bodies and financing instruments in general, may not be entirely immune to potential biases and conflicts of interests. A further point is that criticality as a concept seems to align with the STA, 2024i’s () notion of foundational technologies without which the Internet may function poorly and the creation of new software may become difficult. The frequent funding of OSS supply chains further aligns with the OpenSSF, 2024e’s () emphasis of dependencies for defining criticality. However, there is also some room for criticism regarding the concept of criticality and its relation to the funding decisions. For instance, STA, 2024g has granted $450$ thousand euros for a project seeking to improve documentation about a JavaScript library. While documentation is important, it remains debatable how critical the library is with respect to the foundational technologies emphasized by the STA.
@@
Turning to RQ.2, the monetary amounts granted are visualized in Fig. 2. (It can be noted that at the time of writing the exchange rate between the EU’s euro and the United States dollar was almost even, and thus no conversion was needed.) The average is a little below four thousand euros, but there are a few outlying projects that have received over a million euros. Without prior work, it is rather difficult to contemplate how large or small these amounts are. Nevertheless, consider an imaginary example of a professional software engineer earning roughly around five thousand euros a month; then, the average funding amount would already account for about six full-time software engineers working with a project for a year. This example allows to argue that the funding amounts are decent enough in the OSS domain. Although the OpenSSF does not provide data on the funding durations, most (84%) of the STA’s funding decisions have been for two years. This observation further reinforces the point about the interpretation of the money involved.
@@
The types of cyber security funded (RQ.4) summarized in Table 2 expectedly align closely with the earlier thematic summary in Table 1. In other words, supply chain security has been the most typical thematic cyber security area funded. In this regard, it is important to further note that some of the OSS projects funded have also addressed the design and implementation of so-called software bills of materials (SBOMs). These materials for keeping track of dependencies and other components are noteworthy for two reasons, the first being that SBOMs are also under active research in software engineering (Bi et al.,, 2024; Xia et al.,, 2023, among others). The second reason is that SBOMs too are present in the EU’s CRA, which mandates commercial software vendors to implement SBOMs for their products and deliver these to market surveillance authorities upon request (Ruohonen and Timmers,, 2024). Although OSS projects and their so-called stewards are excluded from this regulatory requirement, the integration of OSS components into commercial software products might partially explain also the funding decisions for SBOMs by the STA and OpenSSF. Another partial explanation might be that also the funding bodies perceive SBOMs as important for improving supply chain security and software security of OSS in general.
Then, regarding RQ.5, already the summary in Table 2 is enough to conclude that there is a divergence between the funding bodies also with respect to the types of cyber security funded. In particular, the STA has funded also many projects addressing encryption and cryptography, network security, and security testing, whereas the OpenSSF has been particularly active in funding of security audits, including those done by third parties. Audits and auditing are important to underline because these have involved also commercial cyber security companies doing the auditing. Thus, the money involved may sometimes go to the private sector even though the goal of funding remains the same. Furthermore, audits and auditing align with the goals of the CRA regulation.
@@
There are also some projects that have received funding from both OSS funding bodies. The examples include FreeBSD and the Python Software Foundation, which is maintaining the popular and large repository for distributing software packages written in Python. The data is again too limited for far-reaching speculations, but, nevertheless, an argument can be raised that the two funding bodies might benefit from joint coordination. This point can be strengthened by noting that the data contains also some cases indicating joint ventures between different funds (OpenSSF, 2024d, ). Against this backdrop, it might be even reasonable to argue that the whole OSS funding landscape might benefit from a high-level coordination body through which all funds could collaborate.
@@
With respect to RQ.3, the funding rationale themes are summarized in Table 3. To put aside cyber security itself, which has sometimes been justified with explicit mentions of critical infrastructure (STA, 2024b, ), also the sustainability theme is strongly present as a rationale. Sustainability as a theme also aligns with the rather frequent rationales of software testing and plain old bug fixing. However, it is worthwhile to mention that some projects have approached the sustainability conundrum through identifying and then deprecating components and dependencies that have reached their end-of-life software life cycle stages or became technologically risky (STA, 2024e, ; STA, 2024f, ). In a similar vein, the earlier point about technological debt is visible in the few projects that have approached sustainability problems through wholesale rewrites and re-implementations. Such solutions are often interpreted as a sign of technical debt of the most severe kind (Verdecchia et al.,, 2021). However, the rewrites and re-implementations have not been only about technological debt and sustainability in general.
@@
The reason is that some of these have involved also transformations to other programming languages; hence, the rewrites and re-implementations also align with the memory safety theme in the earlier Table 2. The Rust programming language is the prime example in this regard. The language and its ecosystem have also received funding from both bodies. From the STA, 2024h the Rust project received funding for implementing SBOMs, while from the OpenSSF, 2024c it received funding for a myriad of security themes, including hiring a dedicated security engineer, audits, documentation, advocacy, and education, as well as new security tools and an establishment of a yet another grant program for mentoring and education. The Rust project’s own grant program yet again highlights the potential benefits from better coordination. The OpenSSF, 2024c’s () funding case also indicates that the funding applications and their evaluations might benefit from a more structured approach. For instance, like in large-scale scientific grant applications, it could be contemplated whether structuring the themes and rationales into work packages would bring more structure and coherence. A further point is that the education, outreach, and documentation involved, and the accompanying overall theme in Table 3, again frames the OSS funding topic also toward academia. In this regard, it could be argued that the security themes funded by the STA and OpenSSF are also something that should be taught in higher education. The Rust case also demonstrates that there might be a need in the nearby future to update curricula with respect to programming languages.
@@
Finally, it is important to emphasize that the funding rationales have involved also implementing new features explicitly or implicitly related to software security and cyber security in general. In addition, also many other general software engineering themes are present, including performance and scalability, interoperability, reliability, usability, and software engineering processes and practices. These rationales allow to end the qualitative exposition by arguing that the tragedy of the (digital) commons alone can neither explain nor elaborate the whole topic of OSS funding. Also, many other explanations are presumably present. The PPP aspect noted is among these explanations; the private sector also greatly benefits from OSS. Thus, further research is needed to better understand the business aspects behind OSS funding. Concepts such as so-called coopetition (Duc et al.,, 2017) might provide a theoretical path to follow in this regard.
6 Discussion
6.1 Conclusion and Contributions
@@
However, cyber security and sustainability alone cannot fully explain the rationales for funding; therefore, neither is the tragedy of the (digital) commons sufficient as an overall theoretical explanation. In addition to raising this important argument, the paper made a notable contribution by linking together the research branches addressing the sustainability of OSS projects and critical infrastructure. Regarding the latter research branch, software’s role has often been downplayed to some extent, and the role of OSS has been almost nonexistent in existing research. A further contribution was made by elaborating how the OSS cyber security funding theme examined is related to regulations. Particularly the EU’s new CRA regulation is relevant also for open source software projects. Against this backdrop, the paper further contributed to recent efforts to better understand the EU’s new funding instruments (Ruohonen and Timmers,, 2025). It seems that the two funding bodies examined have patched either intentional or unintentional omissions in more general public sector funding in Europe and likely also elsewhere. In a similar vein, also threat intelligence reports about supply chain security have missed the sustainability of OSS projects as a mitigative factor (Hamer et al.,, 2025). Finally, the paper contributed to the still nascent theorization and framing of funding for open source software projects.
The point about theorization and framing deserves three additional brief remarks. The first remark is that the involvement of the private sector aligns the OSS funding topic toward so-called public-private partnerships (PPPs), which have been common in the cyber security domain in general (Ruohonen,, 2020). Insofar as OSS is interpreted to be on the public side, which would seem a reasonable interpretation with respect to the concept of public goods underneath the tragedy of the commons, PPPs are particularly visible in the OpenSSF case because the associated fund was established by donations from large technology companies, among them Amazon, Google, and Microsoft (OpenSSF,, 2023). The second remark follows: PPPs are typically also required for funding of large-scale applied scientific projects, including those funded by the EU’s Horizon Europe program. These two remarks logically lead to the final remark: the OSS funding topic is also about the so-called technology transfer model, which has been endorsed and debated in software engineering research for a long time (Mikkonen et al.,, 2018; Wohlin et al.,, 2024, among many others). Regarding the transfers involved, however, scientific research and academia are missing. Given the Germany’s involvement behind the STA, the transfers are primarily between the public sector, the public (OSS) domain, and the private sector.
Therefore, it could be contemplated whether adding academic software engineering research to the mixture would improve the security of OSS even further, perhaps opening also new collaborative avenues and further funding possibilities. This contemplation is reinforced by the educational efforts done by some OSS projects with the help of funding from the two bodies. There are also some indications in the data that some projects have already justified their funding applications with references to scientific research projects and the EU’s own software development projects (STA, 2024e, ). Thus, the earlier points about coordination can be extended toward a larger governance space, which, in Europe, involves also the cyber security regulations mentioned.
6.2 Further Work
@@
The concept of criticality is the keyword behind the funding decisions. However, as was discussed, it is not entirely clear how the two funding bodies have used the concept in decision-making about funding applications. For instance, the corresponding working group established by OpenSSF, 2024e has emphasized dependencies, popularity, and related concepts in determining a project’s criticality. However, dependencies and popularity may align poorly with the STA’s notions about open digital infrastructures and foundational technologies. Consider DNS and BGP implementations as an example; they are clearly critical for the Internet’s functioning, but they may not be as popular as some other OSS projects with large user bases. Nor do they garner a large amount of dependencies. Thus, further research is required about the criticality concept and its practical use in funding decisions.
@@
In addition, an important point regarding further research is the evaluation of not only funding applications but the successes or failures of the projects funded. In this regard, OpenSSF, (2023) has noted that their evaluation has been “largely anecdotal” and that further goals include “highlighting the success factors to future grant recipients and using them in our grant review process”. The STA, 2024a’s () evaluation criteria for funding applications are a little more systematic; prevalence, relevance, sustainability, public interests, activities planned and their feasibility, and expertise of the applicants are evaluated. However, the Agency’s online materials do not say anything about whether funded OSS projects are evaluated post hoc after their completions. The evaluation criteria and related points could be investigated in further research by means of interviews, focus groups, or even surveys, given the apparent proliferation of different funds for OSS projects.
@@
Different evaluation criteria have been considered and compared also in existing research (Osborne et al.,, 2024), but it remains unclear how applicable they are in practice. Science funding provides a good example. On one hand, learning from the funding of scientific projects might help with the goal of improving evaluations. Provided that the point about potential biases and conflicts of interests is true even to some extent, the learning might offer means to improve objectivity in reviewing because objectivity has traditionally been highly valued in science. For instance, it is unclear whether the STA and OpenSSF have used external reviewers, and whether such reviewers would be seen as preferable. On the other hand, it could be also argued that science funding might be a poor reference because it is often rather bureaucratic.